# Logarithmic regret in the ergodic Avellaneda–Stoikov market making model

### v1

@@ A.4. Proof of Theorem 5
By using the comparison principle (see [27, Proposition 3.2]), we know that $\varphi^{c_{1}}(t,q)\leq\mu(t,q)$ for $(t,q)\in[0,+\infty)\times\Omega^{Q}$. We then consider $c_{2}=\sup_{q\in\Omega^{Q}}\big{(}G(q)-\hat{v}(q)\big{)}$, and clearly $\varphi^{c_{2}}(0,q)\geq\mu(0,q)$. By using the comparison principle again, we have $\varphi^{c_{2}}(t,q)\geq\mu(t,q)$. Therefore,
@@ A.7. Proof of Proposition 7
To prove Proposition 7, we recommend to follow the idea in [27, Proposition 4.7] and use the properties of the Hamiltonian function $H$ in Lemma 35. ∎
@@ A.9. Proof of Theorem 9
$\begin{split}0=-\phi q^{2}-\gamma&+\frac{\lambda^{+}}{\kappa}\exp{\big{(}-1-\kappa\hat{v}(q)+\kappa\hat{v}(q-1)\big{)}}\bm{1}_{q>\underline{q}}\\ &+\frac{\lambda^{-}}{\kappa}\exp{\big{(}-1-\kappa\hat{v}(q)+\kappa\hat{v}(q+1)\big{)}}\bm{1}_{q<\bar{q}}\,.\end{split}$
@@
(39) $-\kappa(\phi q^{2}+\gamma)+\lambda^{+}e^{-1}\frac{\hat{\omega}(q-1)}{\hat{\omega}(q)}\bm{1}_{q>\underline{q}}+\lambda^{-}e^{-1}\frac{\hat{\omega}(q+1)}{\hat{\omega}(q)}\bm{1}_{q<\bar{q}}=0\,.$
@@
$\bm{C}=\begin{bmatrix}-\kappa(\phi\bar{q}^{2}+\gamma)&\lambda^{+}e^{-1}&0&&&...&\\ \lambda^{-}e^{-1}&-\kappa\big{(}\phi(\bar{q}-1)^{2}+\gamma\big{)}&\lambda^{+}e^{-1}&&&...&\\ &&&...&\\ &...&&&\lambda^{-}e^{-1}&-\kappa\big{(}\phi(\underline{q}+1)^{2}+\gamma\big{)}&\lambda^{+}e^{-1}\\ &...&&&0&\lambda^{-}e^{-1}&-\kappa(\phi\underline{q}^{2}+\gamma)\end{bmatrix}$
@@
By the fact of $\gamma=\frac{\lambda_{max}}{\kappa}$, we observe that $\bm{C}=\bm{A}-\lambda_{max}\bm{I}$, where the matrix $\bm{A}$ is given in Theorem 2, $\lambda_{max}$ is the largest eigenvalue of $\bm{A}$ and $\bm{I}$ is the identity matrix. As $\bm{A}$ has $n$ distinct eigenvalues as proved in Appendix A.5, hence $\text{rank}(\bm{C})=n-1$. Therefore, $\bm{\hat{\omega}}$ is the set of non-trivial solutions to the homogeneous equation (40) with the coefficient matrix $\bm{C}$. ∎
@@ A.10. Proof of Lemma 11
Given the ergodic optimal control $\psi(q),\,\forall q\in\Omega^{Q}$, we consider the dynamic of inventory process (4) as a discrete Markov chain. Let $\bm{P}$ be an $n-$square matrix labelled from $\bar{q}$ to $\underline{q}$ denoting the state transition matrix. The entries of $\bm{P}$ are
(41) $\bm{P}_{ij}=\begin{cases}1-\lambda^{+}e^{-\kappa\psi^{+}(i)},&\text{if $i=j=\bar{q}$},\\ 1-\lambda^{-}e^{-\kappa\psi^{-}(i)},&\text{if $i=j=\underline{q}$},\\ 1-\lambda^{+}e^{-\kappa\psi^{+}(i)}-\lambda^{-}e^{-\kappa\psi^{-}(i)}+2\lambda^{+}\lambda^{-}e^{-\kappa(\psi^{+}(i)+\psi^{-}(i))},&\text{if $i=j\neq\bar{q},\underline{q}$},\\ \lambda^{-}e^{-\kappa\psi^{-}(i)}(1-\lambda^{+}e^{-\kappa\psi^{+}(i)}),&\text{if $i=j-1\neq\underline{q}$},\\ \lambda^{+}e^{-\kappa\psi^{+}(i)}(1-\lambda^{-}e^{-\kappa\psi^{-}(i)}),&\text{if $i=j+1\neq\bar{q}$},\\ \lambda^{-}e^{-\kappa\psi^{-}(i)},&\text{if $i=j-1=\underline{q}$},\\ \lambda^{+}e^{-\kappa\psi^{+}(i)},&\text{if $i=j+1=\bar{q}$},\\ 0,&\text{otherwise}.\end{cases}$
From the discussion for the tridiagonal matrix in Appendix A.5, clearly $\bm{P}$ is diagonalisable. Let $\lambda_{1},\lambda_{2},...,\lambda_{n}$ be $n$ real eigenvalues of $\bm{P}$. Let $\bm{\Lambda}$ be the diagonal matrix of $\bm{P}$ with $\bm{\Lambda}=\bm{U}^{-1}\bm{P}\bm{U}$, where $\bm{U}$’s columns are the eigenvectors of $\bm{P}$. As $\bm{P}$ is a stochastic matrix, hence by Perron–Frobenius theorem,
$1=\lambda_{1}>|\lambda_{2}|>...>|\lambda_{n}|.$
Therefore, given an initial distribution $\bm{q}$, the equilibrium $\bm{\pi}$ under $\delta^{*,\pm}$ as $T$ goes to infinity is
$\begin{split}\bm{\pi}&=\lim_{T\rightarrow+\infty}\bm{q}\bm{P}^{T}\\ &=\bm{q}\lim_{T\rightarrow+\infty}\bm{U}\bm{\Lambda}^{T}\bm{U}^{-1}\\ &=\bm{q}\bm{U}\bm{W}\bm{U}^{-1}\,,\end{split}$
with $\bm{W}$ an $n-$square matrix with only $1$ on the first diagonal element and 00 otherwise
@@ A.11. Proof of Lemma 17
The equation (24) is a step in the proof of Theorem 27 in a simpler case. We know that $\psi^{\kappa^{\star}}$ is the optimal control for the ergodic market making model under $\kappa^{\star}$ satisfying (18). Hence, from the ergodic HJB equation (16), we have
$\begin{split}0=-\gamma(\kappa^{\star})-\phi q^{2}+&\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa^{\star},+}(q)}\Big{(}\hat{v}^{\kappa^{\star}}(q-1)-\hat{v}^{\kappa^{\star}}(q)+\psi^{\kappa^{\star},+}(q)\Big{)}\bm{1}_{q>\underline{q}}\\ +&\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa^{\star},-}(q)}\Big{(}\hat{v}^{\kappa^{\star}}(q+1)-\hat{v}^{\kappa^{\star}}(q)+\psi^{\kappa^{\star},-}(q)\Big{)}\bm{1}_{q<\bar{q}}\,.\end{split}$
Therefore,
$\begin{split}\gamma(\kappa^{\star})=\lambda^{+}\psi^{\kappa^{\star},+}(q)e^{-\kappa^{\star}\psi^{\kappa^{\star},+}(q)}&+\lambda^{-}\psi^{\kappa^{\star},-}(q)e^{-\kappa^{\star}\psi^{\kappa^{\star},-}(q)}-\phi q^{2}\\ &+\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa^{\star},+}(q)}\big{(}\hat{v}^{\kappa^{\star}}(q-1)-\hat{v}^{\kappa^{\star}}(q)\big{)}\bm{1}_{q>\underline{q}}\\ &+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa^{\star},-}(q)}\big{(}\hat{v}^{\kappa^{\star}}(q+1)-\hat{v}^{\kappa^{\star}}(q)\big{)}\bm{1}_{q<\bar{q}}\,,\end{split}$
where we ignore the indicator functions in the first line and use $\psi^{\kappa^{\star},\pm}(q)e^{-\kappa^{\star}\psi^{\kappa^{\star},\pm}(q)}=0$ for $q=\bar{q},\underline{q}$ respectively. Moreover, we notice that the first line satisfies (21), therefore
$\begin{split}\bigg{|}\gamma(\kappa^{\star})T-\mathbb{E}_{q}\Big{[}&\int_{0}^{T}f(Q_{t}^{\psi^{\kappa^{\star}};\kappa^{\star}},\psi^{\kappa^{\star}};\kappa^{\star})\,\mathrm{d}t\Big{]}\bigg{|}\\ &=\bigg{|}\mathbb{E}_{q}\Big{[}\int_{0}^{T}\big{(}\gamma(\kappa^{\star})-f(Q_{t}^{\psi^{\kappa^{\star}};\kappa^{\star}},\psi^{\kappa^{\star}};\kappa^{\star})\big{)}\,\mathrm{d}t\Big{]}\bigg{|}\\ &=\bigg{|}\mathbb{E}_{q}\Big{[}\int_{0}^{T}\Big{(}\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa^{\star},+}(q)}\big{(}\hat{v}^{\kappa^{\star}}(q-1)-\hat{v}^{\kappa^{\star}}(q)\big{)}\bm{1}_{q>\underline{q}}\\ &\qquad\qquad\qquad\quad+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa^{\star},-}(q)}\big{(}\hat{v}^{\kappa^{\star}}(q+1)-\hat{v}^{\kappa^{\star}}(q)\big{)}\bm{1}_{q<\bar{q}}\Big{)}\,\mathrm{d}t\Big{]}\bigg{|}\,.\end{split}$
Let $\pi^{\kappa^{\star}}$ denote the equilibrium of the optimal ergodic market making model under parameter $\kappa^{\star}$ and function $h$ be
$\begin{split}h(\kappa^{\star},q)=\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa^{\star},+}(q)}\big{(}\hat{v}^{\kappa^{\star}}(q-1)&-\hat{v}^{\kappa^{\star}}(q)\big{)}\bm{1}_{q>\underline{q}}\\ &+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa^{\star},-}(q)}\big{(}\hat{v}^{\kappa^{\star}}(q+1)-\hat{v}^{\kappa^{\star}}(q)\big{)}\bm{1}_{q<\bar{q}}\,.\end{split}$
By Lemma 36 (2), $h$ is bounded by $\bar{h}\in\mathbb{R}^{+}$. From a simpler version of Proposition 26 by substituting $\psi^{\kappa}$ to $\psi^{\kappa^{\star}}$, Lemma 12 and Lemma 36 (2), we have
$\begin{split}\bigg{|}\gamma(\kappa^{\star})T-\mathbb{E}_{q}\Big{[}&\int_{0}^{T}f(Q_{t}^{\psi^{\kappa^{\star}};\kappa^{\star}},\psi^{\kappa^{\star}};\kappa^{\star})\,\mathrm{d}t\Big{]}\bigg{|}\\ &=\bigg{|}\int_{0}^{T}\int_{\Omega^{Q}}\Big{(}\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa^{\star},+}(q)}\big{(}\hat{v}^{\kappa^{\star}}(q-1)-\hat{v}^{\kappa^{\star}}(q)\big{)}\bm{1}_{q>\underline{q}}\\ &\qquad\qquad\qquad\quad+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa^{\star},-}(q)}\big{(}\hat{v}^{\kappa^{\star}}(q+1)-\hat{v}^{\kappa^{\star}}(q)\big{)}\bm{1}_{q<\bar{q}}\Big{)}\,\mathrm{d}t\,\mathrm{d}\pi^{\kappa^{\star}}_{t}\bigg{|}\\ &\leq\bigg{|}\bar{h}\int_{0}^{T}\int_{\Omega^{Q}}\frac{h(\kappa^{\star},q)}{\bar{h}}(\mathrm{d}\pi^{\kappa^{\star}}_{t}-\mathrm{d}\pi^{\kappa^{\star}})\,\mathrm{d}t\bigg{|}\\ &\leq\bar{h}\int_{0}^{T}\left\|\pi^{\kappa^{\star}}_{t}-\pi^{\kappa^{\star}}\right\|_{TV}\,\mathrm{d}t\leq\frac{\bar{h}C}{-\ln\alpha}\,,\end{split}$
A.12. Proof of Lemma 20
Proof.
@@ A.12. Proof of Lemma 20
$0=\partial_{t}w(t,q)-rw(t,q)+H\Big{(}q,\psi^{\kappa},(w(t,q^{\prime})-w(t,q))_{q^{\prime}\in\{q-1,q+1\}};\kappa^{\star}\Big{)},\,\forall q\in\Omega^{Q},$
@@
Let us consider the following stochastic process, and we omit the superscript for $Q_{t}^{\psi^{\kappa};\kappa^{\star}}$ for notational simplicity.
$X(s)=e^{-r(s-t)}w(s,Q_{s})+\int_{t}^{s}e^{-r(u-t)}f(Q_{u},\psi^{\kappa}_{u};\kappa^{\star})\,\mathrm{d}u$
We know that $Q_{t}^{\psi^{\kappa};\kappa^{\star}}$ follows the SDE (4) with market parameter $\kappa^{\star}$ and control $\psi^{\kappa}$, i.e.
$\begin{split}dQ_{t}^{\psi^{\kappa};\kappa^{\star}}&=dN_{t}^{\psi^{\kappa},-}-dN_{t}^{\psi^{\kappa},+}\\ &=\big{(}\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,-}}-\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,+},}\big{)}dt+d\tilde{N}_{t}^{\psi^{\kappa},-}-d\tilde{N}_{t}^{\psi^{\kappa},+},\end{split}$
@@
$dX(s)=\,e^{-r(s-t)}\Big{\{}\partial_{s}w(s,Q_{s})-rw(s,Q_{s})+\lambda^{+}e^{-\kappa^{\star}\psi^{\kappa,+}}\left(w(s,Q_{s}-1)-w(s,Q_{s})\right)\bm{1}_{Q_{s}>\underline{q}}$
$+\lambda^{-}e^{-\kappa^{\star}\psi^{\kappa,-}}\left(w(s,Q_{s}+1)-w(s,Q_{s})\right)\bm{1}_{Q_{s}<\bar{q}}+f(Q_{s},\psi^{\kappa};\kappa^{\star})\Big{\}}ds$
$+e^{-r(s-t)}\Big{\{}\left(w(s,Q_{s}-1)-w(s,Q_{s})\right)\bm{1}_{Q_{s}>\underline{q}}d\tilde{N}_{s}^{+}+\left(w(s,Q_{s}+1)-w(s,Q_{s})\right)\bm{1}_{Q_{s}<\bar{q}}d\tilde{N}_{s}^{-}\Big{\}}$
@@
$+e^{-r(s-t)}\Big{\{}\left(w(s,Q_{s}-1)-w(s,Q_{s})\right)\bm{1}_{Q_{s}>\underline{q}}d\tilde{N}_{s}^{+}+\left(w(s,Q_{s}+1)-w(s,Q_{s})\right)\bm{1}_{Q_{s}<\bar{q}}d\tilde{N}_{s}^{-}\Big{\}}.$
where the last equality comes from the definition of $H(\cdot;\kappa^{\star})$ (27) and $f(\cdot;\kappa^{\star})$ (21). Take the integral and expectation for $X(s)$, we have
@@
$\begin{split}w(t,q)=\mathbb{E}\big{[}X(t)\big{|}Q_{t}=q\big{]}&=\mathbb{E}\big{[}X(T)\big{|}Q_{t}=q\big{]}\\ &=\mathbb{E}_{q}\Big{[}\int_{t}^{T}e^{-r(u-t)}f(Q_{u},\psi^{\kappa};\kappa^{\star})\,\mathrm{d}u+e^{-r(T-t)}G(Q_{T})\Big{]}.\end{split}$
Hence $w(t,q)=v_{r}^{\psi^{\kappa}}(t,q;T;\kappa^{\star})$ by (28). ∎
A.13. Proof of Proposition 21

### v2

@@ A.4. Proof of Theorem 5
By using the comparison principle (see [26, Proposition 3.2]), we know that $\varphi^{c_{1}}(t,q)\leq\mu(t,q)$ for $(t,q)\in[0,+\infty)\times\Omega^{Q}$. We then consider $c_{2}=\sup_{q\in\Omega^{Q}}\big{(}G(q)-\hat{v}(q)\big{)}$, and clearly $\varphi^{c_{2}}(0,q)\geq\mu(0,q)$. By using the comparison principle again, we have $\varphi^{c_{2}}(t,q)\geq\mu(t,q)$. Therefore,
@@ A.7. Proof of Proposition 7
To prove Proposition 7, we recommend to follow the idea in [26, Proposition 4.7] and use the properties of the Hamiltonian function $H$ in Lemma 37. ∎
@@ A.9. Proof of Theorem 9
$\begin{split}0=-\phi q^{2}-\gamma&+\frac{\lambda^{+}}{\kappa}\exp{\big{(}-1-\kappa\hat{v}(q)+\kappa\hat{v}(q-1)\big{)}}\mathds{1}_{q>\underline{q}}\\ &+\frac{\lambda^{-}}{\kappa}\exp{\big{(}-1-\kappa\hat{v}(q)+\kappa\hat{v}(q+1)\big{)}}\mathds{1}_{q<\bar{q}}\,.\end{split}$
@@
(43) $-\kappa(\phi q^{2}+\gamma)+\lambda^{+}e^{-1}\frac{\hat{\omega}(q-1)}{\hat{\omega}(q)}\mathds{1}_{q>\underline{q}}+\lambda^{-}e^{-1}\frac{\hat{\omega}(q+1)}{\hat{\omega}(q)}\mathds{1}_{q<\bar{q}}=0\,.$
@@
$\bm{C}=\begin{bmatrix}-\kappa(\phi\bar{q}^{2}+\gamma)&\lambda^{+}e^{-1}&0&&&...&\\ \lambda^{-}e^{-1}&-\kappa\big{(}\phi(\bar{q}-1)^{2}+\gamma\big{)}&\lambda^{+}e^{-1}&&&...&\\ &&&...&\\ &...&&&\lambda^{-}e^{-1}&-\kappa\big{(}\phi(\underline{q}+1)^{2}+\gamma\big{)}&\lambda^{+}e^{-1}\\ &...&&&0&\lambda^{-}e^{-1}&-\kappa(\phi\underline{q}^{2}+\gamma)\end{bmatrix}\,.$
@@
Due to the fact of $\gamma=\frac{\lambda_{max}}{\kappa}$, we observe that $\bm{C}=\bm{A}-\lambda_{max}\bm{I}$, where the matrix $\bm{A}$ is given in Theorem 2, $\lambda_{max}$ is the largest eigenvalue of $\bm{A}$ and $\bm{I}$ is the identity matrix. As $\bm{A}$ has $n$ distinct eigenvalues as proved in Appendix A.5, hence $\text{rank}(\bm{C})=n-1$. Therefore, the null space of the matrix $\bm{C}$ has dimension $1$ by rank-nullity theorem, implying that the solution $\bm{\hat{\omega}}$ to the homogeneous equation (44) is unique (up to multiplicative factors). Indeed, $\bm{\hat{\omega}}$ is the eigenvector corresponding the dominant eigenvalue of matrix $A$, which is a Metzler matrix (non-negative off-diagonal entries). By Perron–Frobenius theorem, the eigenvector to the dominant eigenvalue is positive, which completes the proof. ∎
@@ A.10. Proof of Lemma 14
The equation (25) is a step in the proof of Theorem 30 in a simpler case. We know that $\psi^{\kappa^{\ast}}$ is the optimal control for the ergodic market making model under $\kappa^{\ast}$ satisfying (19). Hence, from the ergodic HJB equation (16), we have
$\begin{split}0=-\gamma(\kappa^{\ast})-\phi q^{2}+&\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa^{\ast},+}(q)}\Big{(}\hat{v}^{\kappa^{\ast}}(q-1)-\hat{v}^{\kappa^{\ast}}(q)+\psi^{\kappa^{\ast},+}(q)\Big{)}\mathds{1}_{q>\underline{q}}\\ +&\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa^{\ast},-}(q)}\Big{(}\hat{v}^{\kappa^{\ast}}(q+1)-\hat{v}^{\kappa^{\ast}}(q)+\psi^{\kappa^{\ast},-}(q)\Big{)}\mathds{1}_{q<\bar{q}}\,.\end{split}$
Therefore,
$\begin{split}\gamma(\kappa^{\ast})=\lambda^{+}\psi^{\kappa^{\ast},+}(q)e^{-\kappa^{\ast}\psi^{\kappa^{\ast},+}(q)}&+\lambda^{-}\psi^{\kappa^{\ast},-}(q)e^{-\kappa^{\ast}\psi^{\kappa^{\ast},-}(q)}-\phi q^{2}\\ &+\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa^{\ast},+}(q)}\big{(}\hat{v}^{\kappa^{\ast}}(q-1)-\hat{v}^{\kappa^{\ast}}(q)\big{)}\mathds{1}_{q>\underline{q}}\\ &+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa^{\ast},-}(q)}\big{(}\hat{v}^{\kappa^{\ast}}(q+1)-\hat{v}^{\kappa^{\ast}}(q)\big{)}\mathds{1}_{q<\bar{q}}\,,\end{split}$
where we ignore the indicator functions in the first line and use $\psi^{\kappa^{\ast},\pm}(q)e^{-\kappa^{\ast}\psi^{\kappa^{\ast},\pm}(q)}=0$ for $q=\bar{q},\underline{q}$ respectively. Moreover, we notice that the first line satisfies (22), therefore
$\begin{split}\bigg{|}\gamma(\kappa^{\ast})T-\mathbb{E}_{q}\Big{[}&\int_{0}^{T}f(Q_{t}^{\psi^{\kappa^{\ast}};\kappa^{\ast}},\psi^{\kappa^{\ast}};\kappa^{\ast})\,\mathrm{d}t\Big{]}\bigg{|}\\ &=\bigg{|}\mathbb{E}_{q}\Big{[}\int_{0}^{T}\big{(}\gamma(\kappa^{\ast})-f(Q_{t}^{\psi^{\kappa^{\ast}};\kappa^{\ast}},\psi^{\kappa^{\ast}};\kappa^{\ast})\big{)}\,\mathrm{d}t\Big{]}\bigg{|}\\ &=\bigg{|}\mathbb{E}_{q}\Big{[}\int_{0}^{T}\Big{(}\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa^{\ast},+}(q)}\big{(}\hat{v}^{\kappa^{\ast}}(q-1)-\hat{v}^{\kappa^{\ast}}(q)\big{)}\mathds{1}_{q>\underline{q}}\\ &\qquad\qquad\qquad\quad+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa^{\ast},-}(q)}\big{(}\hat{v}^{\kappa^{\ast}}(q+1)-\hat{v}^{\kappa^{\ast}}(q)\big{)}\mathds{1}_{q<\bar{q}}\Big{)}\,\mathrm{d}t\Big{]}\bigg{|}\,.\end{split}$
Let $\pi^{\kappa^{\ast}}$ denote the equilibrium of the optimal ergodic market making model under parameter $\kappa^{\ast}$ and function $h$ be
$\begin{split}h(\kappa^{\ast},q)=\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa^{\ast},+}(q)}\big{(}\hat{v}^{\kappa^{\ast}}(q-1)&-\hat{v}^{\kappa^{\ast}}(q)\big{)}\mathds{1}_{q>\underline{q}}\\ &+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa^{\ast},-}(q)}\big{(}\hat{v}^{\kappa^{\ast}}(q+1)-\hat{v}^{\kappa^{\ast}}(q)\big{)}\mathds{1}_{q<\bar{q}}\,.\end{split}$
By Lemma 38 (2), $h$ is bounded by $\bar{h}\in\mathbb{R}^{+}$. From a simpler version of Proposition 29 by substituting $\psi^{\kappa}$ to $\psi^{\kappa^{\ast}}$, Lemma 28 and Lemma 38 (2), we have
$\begin{split}\bigg{|}\gamma(\kappa^{\ast})T-\mathbb{E}_{q}\Big{[}&\int_{0}^{T}f(Q_{t}^{\psi^{\kappa^{\ast}};\kappa^{\ast}},\psi^{\kappa^{\ast}};\kappa^{\ast})\,\mathrm{d}t\Big{]}\bigg{|}\\ &=\bigg{|}\int_{0}^{T}\int_{\Omega^{Q}}\Big{(}\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa^{\ast},+}(q)}\big{(}\hat{v}^{\kappa^{\ast}}(q-1)-\hat{v}^{\kappa^{\ast}}(q)\big{)}\mathds{1}_{q>\underline{q}}\\ &\qquad\qquad\qquad\quad+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa^{\ast},-}(q)}\big{(}\hat{v}^{\kappa^{\ast}}(q+1)-\hat{v}^{\kappa^{\ast}}(q)\big{)}\mathds{1}_{q<\bar{q}}\Big{)}\,\mathrm{d}t\,\mathrm{d}\pi^{\kappa^{\ast}}_{t}\bigg{|}\\ &\leq\bigg{|}\bar{h}\int_{0}^{T}\int_{\Omega^{Q}}\frac{h(\kappa^{\ast},q)}{\bar{h}}(\mathrm{d}\pi^{\kappa^{\ast}}_{t}-\mathrm{d}\pi^{\kappa^{\ast}})\,\mathrm{d}t\bigg{|}\\ &\leq\bar{h}\int_{0}^{T}\left\|\pi^{\kappa^{\ast}}_{t}-\pi^{\kappa^{\ast}}\right\|_{TV}\,\mathrm{d}t\leq\frac{\bar{h}C}{-\ln\alpha}\,,\end{split}$
@@ A.11. Proof of Lemma 18
$0=\partial_{t}w(t,q)-rw(t,q)+H\Big{(}q,\psi^{\kappa},(w(t,q^{\prime})-w(t,q))_{q^{\prime}\in\{q-1,q+1\}};\kappa^{\ast}\Big{)},\,\forall q\in\Omega^{Q},$
@@
Let us consider the following stochastic process, and we omit the superscript for $Q_{t}^{\psi^{\kappa};\kappa^{\ast}}$ for notational simplicity.
$X(s)=e^{-r(s-t)}w(s,Q_{s})+\int_{t}^{s}e^{-r(u-t)}f(Q_{u},\psi^{\kappa}_{u};\kappa^{\ast})\,\mathrm{d}u$
We know that $Q_{t}^{\psi^{\kappa};\kappa^{\ast}}$ follows the SDE (4) with market parameter $\kappa^{\ast}$ and control $\psi^{\kappa}$, i.e.
$\begin{split}dQ_{t}^{\psi^{\kappa};\kappa^{\ast}}&=dN_{t}^{\psi^{\kappa},-}-dN_{t}^{\psi^{\kappa},+}\\ &=\big{(}\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,-}}-\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,+},}\big{)}dt+d\tilde{N}_{t}^{\psi^{\kappa},-}-d\tilde{N}_{t}^{\psi^{\kappa},+},\end{split}$
@@
$dX(s)=\,e^{-r(s-t)}\Big{\{}\partial_{s}w(s,Q_{s})-rw(s,Q_{s})+\lambda^{+}e^{-\kappa^{\ast}\psi^{\kappa,+}}\left(w(s,Q_{s}-1)-w(s,Q_{s})\right)\mathds{1}_{Q_{s}>\underline{q}}$
$+\lambda^{-}e^{-\kappa^{\ast}\psi^{\kappa,-}}\left(w(s,Q_{s}+1)-w(s,Q_{s})\right)\mathds{1}_{Q_{s}<\bar{q}}+f(Q_{s},\psi^{\kappa};\kappa^{\ast})\Big{\}}ds$
$+e^{-r(s-t)}\Big{\{}\left(w(s,Q_{s}-1)-w(s,Q_{s})\right)\mathds{1}_{Q_{s}>\underline{q}}d\tilde{N}_{s}^{+}+\left(w(s,Q_{s}+1)-w(s,Q_{s})\right)\mathds{1}_{Q_{s}<\bar{q}}d\tilde{N}_{s}^{-}\Big{\}}$
@@
$+e^{-r(s-t)}\Big{\{}\left(w(s,Q_{s}-1)-w(s,Q_{s})\right)\mathds{1}_{Q_{s}>\underline{q}}d\tilde{N}_{s}^{+}+\left(w(s,Q_{s}+1)-w(s,Q_{s})\right)\mathds{1}_{Q_{s}<\bar{q}}d\tilde{N}_{s}^{-}\Big{\}}.$
where the last equality comes from the definition of $H(\cdot;\kappa^{\ast})$ (31) and $f(\cdot;\kappa^{\ast})$ (22). Take the integral and expectation for $X(s)$, we have
@@
$\begin{split}w(t,q)=\mathbb{E}\big{[}X(t)\big{|}Q_{t}=q\big{]}&=\mathbb{E}\big{[}X(T)\big{|}Q_{t}=q\big{]}\\ &=\mathbb{E}_{q}\Big{[}\int_{t}^{T}e^{-r(u-t)}f(Q_{u},\psi^{\kappa};\kappa^{\ast})\,\mathrm{d}u+e^{-r(T-t)}G(Q_{T})\Big{]}.\end{split}$
Hence $w(t,q)=v_{r}^{\psi^{\kappa}}(t,q;T;\kappa^{\ast})$ by (32). ∎
A.12. Proof of Proposition 19